Directions: a Statement of Comprehensive Income using the following format and calculated values:


The Mirriam Company has the following statement of comprehensive income balances at the end of the year.

Gross Profit = ₱ 2,580,750

Operating Profit = ₱ 1,554,330

Income Tax Expense = 25% of Operating Profit

Comprehensive Income = ₱ 2,000,000


construct a statement of comprehensive income which contains the determination of net income and comprehensive income for the period.



1
Expert's answer
2021-10-05T13:53:59-0400

Gross profit = ₱ 2,580,750

Operating profit = ₱ 1,554,330

Income tax expense = 25% of Operating Profit

Income tax expense = 25% × 1,554,330 = ₱388,582.5

Net Income = Operating profit - income tax expense

=1,554,330 - 388,582.5 = ₱ 1,165,747.5

Comprehensive Income = ₱ 2,000,000

Total Comprehensive Income = 2,000,000 + 1,165,747.5 = ₱ 1,365,747.5




The Mirriam Company

Statement of comprehensive income

at the end of the year

Net income ₱1,165,747

Comprehensive income 2,000,000

Total comprehensive income₱1,365,747